OpenStudy (anonymous):

The sum of four consecutive multiples of 5 is 210. What is the smallest of these numbers and what is the largest?

OpenStudy (tkhunny):

Please name the four numbers.

OpenStudy (anonymous):

if the lowest number of the 4 is denoted by "x" then the next consecutive multiple of 5 can be denoted as "x +5" and so on and so forth or, you can denote the highest number of the 4 as "y" and the previous multiples of 5 would be "y-5n" depends on which way you want to solve it

OpenStudy (anonymous):

FIRST I WOULD LIKE TO KNOW WHAT THE SMALLEST MULTIPLE IS IS. HOW IT IS THE MOST EASY TO SOLVE WOULD BE FINE

OpenStudy (tkhunny):

I'm not particularly happy with that setup, as there is nothing about 'x' that makes it a multiple of 5. It may be fine and it may work, but it is somewhat by luck that this is so. On the other hand, we now have names for the four values. I would prefer: Lowest Number = 5k, with k being a whole number. 2nd Number = 5(k+1) 3rd Number = 5(k+2) Greatest Number = 5(k+3) Now, we find a solution for k and all such questions can be answered.

OpenStudy (anonymous):

THE QUESTION I WAS GIVEN ONLY STATES THAT THE SUM OF FOUR MULTIPLES OF 5 IS 10. IT THEN ASKS WHAT THE SMALLEST MULTIPLE IS?

OpenStudy (tkhunny):

That is correct. Now, solve the problem. The least of the 4 multiples is 5k and the greatest is 5(k+3), assuming there is a Whole Number k that works.

OpenStudy (anonymous):

SORRY.... THE QUESTION STATES THAT THE SUM OF FOUR CONSECUTIVE MULTIPLES OF 5 IS 210. IT THEN ASKS WHAT IS THE SMALLEST MULTIPLE?

OpenStudy (tkhunny):

Are you planning to seek a solution through your own hard work and efforts?

OpenStudy (anonymous):

IF I KNEW HOW TO GET THE ANSWER THAT IS WHAT I WOULD BE DOING BUT SINCE I DO NOT KNOW I AM SEEKING ASSISTANCE FROM SOMEONE WHO DOES SO I CAN LEARN HOW TO ARRIVE AT THE PROPER ANSWER

OpenStudy (tkhunny):

That is not a good response. It is appropriate to assume that you have been given this problem only after you have been given the tools to solve it. Therefore, given enough time and effort, you may be able to arrive at a solution on your own. We have names for the numbers. This was very important. Lowest Number = 5k 2nd Number = 5(k+1) 3rd Number = 5(k+2) Greatest Number = 5(k+3) The problem statement calls for the sum of these four numbers. Can you write an expression that represents the sum of the four number?

OpenStudy (anonymous):

SINCE I REALLY DON'T UNDERSTAND I WILL JUST HAVE TO LOOK FOR SIOME ONE TO GIVE ME A FULL EXAMPLE. THANKS FOR ALL YOUR HELP AND HAVE A BLESSED DAY!

OpenStudy (tkhunny):

Have you considered addition? Given two numbers, say 5, and 7, the sum of these two numbers can be represented as 5+7. That is all that is needed at this time. There are those who will work problem for you, and do your home work for you. Please note that this sort of behavior is not consistent with the Code of Conduct for this website. Note more carefully, though, that such individual CANNOT be with you during your exam. You have shown no effort in solving this problem. This must change or you will fail your course.